Flora, Inc., reported an EPS of $3.3 this year (t0). Flora is expected to maintain a retained earnings ratio of 0.2 and ROE of 0.15 for the next five years. After the fifth year, ROE is expected to decrease to 0.07. Applying the cost of equity of 0.11 and the multi-stage growth model, compute the intrinsic price of Flora.
Answer #1

Growth Rate = Retained Earining Ratio * ROE

= 0.20 * 0.15

= 3% OR 0.03

Growth Rate in second stage = 0.2 * 0.07

= 1.4%  

D1 = 3.3 * .20 = 0.66

D2 = 0.66 ( 1+ 0.03) = 0.6798

D2 = 0.6798 ( 1+ 0.03) = 0.700194

D2 = 0.700194 ( 1+ 0.03) = 0.72119982

D2 = 0.72119982 ( 1+ 0.03) = 0.7428358146

Value of Stock = D D DE + D:(1+G) (1 + r)5 G DA + (1+r)2 + (1 + r)3 + (1+r)4 *(1 + r) 5 (1 + r)

10.66 0.6798 0.700194 0.72119982 0.7428358146 (1+0.11)*(1+0.112*(1+0.11)3(1+0.11) 4 (1+0.11) 0.7428358146(1+0.014) | 011-0.01

= $ 7.23

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.